OpenStudy (anonymous):

how do u work this out √(y^15)

OpenStudy (anonymous):

y^(15/2) e^(i pi floor(1/2-(15 arg(y))/(2 pi)))

OpenStudy (anonymous):

y^(15/2)

OpenStudy (anonymous):

7 remainer 1

OpenStudy (anonymous):

y^(15/2)+O(y^(25/2))

OpenStudy (anonymous):

y^(15/2)+O((1/y)^(25/2))

OpenStudy (anonymous):

d/dy(sqrt(y^15)) = (15 sqrt(y^15))/(2 y)

OpenStudy (anonymous):

where did u get 25

OpenStudy (anonymous):

Possible derivation: d/dy(sqrt(y^15)) | Use the chain rule, d/dy(sqrt(y^15)) = ( du)/( du) ( du)/( dy), where u = sqrt(y^15) and ( du)/( du) = HoldForm'(u): = | HoldForm'(sqrt(y^15)) (d/dy(sqrt(y^15))) | Use the chain rule, d/dy(sqrt(y^15)) = ( dsqrt(u))/( du) ( du)/( dy), where u = y^15 and ( dsqrt(u))/( du) = 1/(2 sqrt(u)): = | HoldForm'(sqrt(y^15)) (d/dy(y^15))/(2 sqrt(y^15)) | The derivative of y^15 is 15 y^14: = | ((15 y^14) HoldForm'(sqrt(y^15)))/(2 sqrt(y^15))

OpenStudy (anonymous):

y = 0

OpenStudy (anonymous):

it say

OpenStudy (lgbasallote):

@best.shakir are you copy-pasting from wolframalpha?

OpenStudy (anonymous):

simplyfy √(y^15)

OpenStudy (anonymous):

y sq root 15

OpenStudy (anonymous):

see here http://www.wolframalpha.com/input/?i=root+of+y%5E15

OpenStudy (anonymous):

yes best shakir l dnt understand

OpenStudy (lgbasallote):

\[\huge \sqrt[m]{x^n}\] does that help @tafara27 ?

OpenStudy (anonymous):

@lgbasallote u are good at explain

OpenStudy (lgbasallote):

uhh wait...wrote that wrong

OpenStudy (lgbasallote):

\[\huge \sqrt[m]{x^n} \implies x^{n/m}\] does that help?

OpenStudy (anonymous):

yea thnks

OpenStudy (lgbasallote):

welcome